Ni(s) + 2Ag+ (1.0M) → Ni2+ (1.0 M) + 2Ag(s) given that Eø for Ni+2/Ni= -0.25, Eø for Ag+/Ag=0.80. Need the emf of the cell.
The Eø should be 0.55 as Eø for Ni+2/Ni= -0.25 but this is an anodic reaction so SOP= -0.25

Grade:12

1 Answers

Aqeel ahmed
4 years ago
E° of cell = 0.80-(-0.25)
               = 0.80+0.25
               =1.05
 Emf of cell =E°-0.059÷n log p÷ R
                   =1.05 - 0.059÷2 ( log 1- 2 log1)
                   =1.05
 Becoz ( log 1- 2 log 1 ) is zero so , product of 0.059 and o is zero....